OpenStudy (anonymous):

How do I solve this problem...? 3/(x^2+3x)+(x+2)/(x+3)<1/x

OpenStudy (anonymous):

\[\frac{ 3 }{ x^2+3x}+\frac{ x+2 }{ x+3 }<\frac{ 1 }{ x }\]

OpenStudy (phi):

I would add -1/x to both sides

OpenStudy (anonymous):

I would get common denomonators

OpenStudy (anonymous):

So after adding -1/x what would the other side look like...?

OpenStudy (phi):

\[ \frac{ 3 }{ x^2+3x}+\frac{ x+2 }{ x+3 }- \frac{ 1 }{ x }<0 \] now do what cc3 suggested. find a common denominator for the 3 fractions it looks like x(x+3)

OpenStudy (anonymous):

So I'd divide it by x+3...? I mean that was my gut feel but then what would happen to the 1/x?

OpenStudy (phi):

no, multiply the 2nd fraction by x/x multiply the last equation by (x+3)/(x+3) now all three fractions will have the same denominator. you can combine the tops

OpenStudy (anonymous):

So that would leave me with \[\frac{ 3 }{ x^2+3x }+\frac{ 2 }{ x^2+3 }-\frac{ x+3 }{ x^2+3 }<0\] ...?

OpenStudy (phi):

the bottoms should be all the same... For example, notice for the last fraction \[ \frac{1}{x}\cdot \frac{(x+3)}{(x+3)} = \frac{(x+3)}{x(x+3)}= \frac{(x+3)}{x^2+3x}\]

OpenStudy (phi):

re do the middle fraction

OpenStudy (anonymous):

OHhhhh so it would actually be \[\frac{ 3 }{ x^2+3 }+\frac{ 2 }{ x^2+3 }+\frac{ x+3 }{ x^2+3 }<0\]

OpenStudy (phi):

no, the minus sign on the third fraction should still be there. the bottom is x^2 + 3x or x(x+3) (which is how I would write it)

OpenStudy (phi):

and the middle fraction is still not correct.

OpenStudy (anonymous):

I'm sorry, could you walk me through this ;3;...? I'm not really good at this kind of thing...

OpenStudy (phi):

\[ \frac{ 3 }{ x^2+3x}+\frac{ x+2 }{ x+3 }- \frac{ 1 }{ x }<0 \] multiply the second fraction by \( \frac{x}{x} \) what do you get? multiply the third fraction by \( \frac{(x+3)}{(x+3)} \)

OpenStudy (anonymous):

\[\frac{ 3 }{ x^2+3x }+\frac{ x^2+2 }{ x^2+3 }-\frac{ x+3 }{ x^2+3 }\]?

OpenStudy (phi):

I guess you forgot how the distributive rule works. you should leave the bottom x(x+3) but if you expand it, you multiply x times *everything inside the parens* you would get x*x + 3*x or x^2 +3x (NOT x^2+3) up top in the second fraction it is x*(x+2) that expands into x*x + 2*x or x^2+2x (NOT x^2 + 2)

OpenStudy (anonymous):

\[\frac{ 3 }{ x^2+3x }+\frac{ x^2+2x }{ x^2+3 }-\frac{ x+3 }{ x^2+3x }\] ?

OpenStudy (anonymous):

\[\frac{ 3 }{ x ^{2}+3 }+\frac{ x+2 }{ x+3 }-\frac{ 1 }{x }<0\] \[or \frac{ 3+x \left( x+2 \right)-1\left( x+3 \right) }{ x \left( x+3 \right) }<0\] \[or~\frac{ 3+x ^{2}+2x-x-3 }{x \left( x+3 \right) }<0\] \[or~\frac{ x \left( x+1 \right) }{ x \left( x+3 \right) }<0\] \[or~\frac{ x+1 }{ x+3 }<0\] either x+1>0,x>-1 and x+3<0,x<-3 or x+1<0,x<-1,x+3>0,x>-3 hence solution is (-3,-1)